OpenStudy (anonymous):

How can you write the expression with a rationalized denominator? (5-√2) / (√3)\

OpenStudy (anonymous):

\[\frac{5-\sqrt2}{\sqrt3}\]?

OpenStudy (anonymous):

multiply top an bottom by \(\sqrt 3\) to get \[\frac{5\sqrt{3}-\sqrt{6}}{3}\]

OpenStudy (anonymous):

& then what? do I have to do that to the numerator?